A company uses the periodic inventory system and had the following activity during the current monthly period.

November 1: Beginning inventory 111 units @ $30
November 5: Purchased 111 units @ $22
November 8: Purchased 61 units @ $30
November 16: Sold 170 units @ $100
November 19: Purchased 70 units @ $30


Using the weighted-average inventory method, the company's ending inventory would be:

Multiple Choice

  • $5,430

  • $5,030

  • $7,260

  • $3,930

  • $3,330
Answer #1

Solution

The correct answer is b) 5,030

  • Units available for Sale = 111 + 111 + 61 + 70 = 353 units
  • Units sold = 170
  • Ending Inventory = 353 – 170 = 183 units
  • Workings

Average Method

Cost of Goods available for sale

Units

Cost/unit

COG for sale

Beginning Inventory

111

$ 30.00

$ 3,330.00

Purchases:

Nov 5

111

$                22.00

$                        2,442.00

Nov 8

61

$ 30.00

$                        1,830.00

Nov 19

70

$ 30.00

$ 2,100.00

TOTAL

353

$              27.49

$ 9,702.00

  • Ending inventory: (353 units − 170 units) * $27.49 = $5,030

Please give thumbs up if you like my answer :-)